A building casts a shadow 26 m long. At the same time, the shadow cast by a 63 cm tall pole is 60 cm long. Find the height of the building.
A. 145.4 m B. 27.3 m C. 25.8 m D. 24.8 m E. 23.3 m

Found 2 solutions by nyc_function, checkley77:
Answer by nyc_function(2741)

I can send a reply to your math question using paint attachment, which is far easier to read and follow.
Write to: nycfunction@yahoo.com
Guido
P.S. Post one question at a time for a faster reply.

Answer by checkley77(12844)

x/26=63/60
60x=63*26
60x=1,638
x=1,638/60
x=27.3 cm tall is the height of the building.
ANSWER. B).